OpenStudy (anonymous):

Write the equation of the line that passes through (–2, 6) and (2, 14) in slope-intercept form.

OpenStudy (anonymous):

can someone help please? medal and fan?

Mehek (mehek14):

slope intercept form: \(\tt{y=\color{blue}{m}x+\color{red}{b}\\ \color{blue}{m=slope} \\ \color{red}{b=y-intercept} }\)

pooja195 (pooja195):

First find the slope \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (anonymous):

To find the slope, use the slope formula which is m= (y2-y1)/(x2-x1). In this case the slope is 2.

OpenStudy (anonymous):

The plug it into the slope intercept formula, which is y = mx+b, so you get 14=2(2) + b, solve for b and you get 3.5 or 7/2.

OpenStudy (anonymous):

so with the 2 what should I do then?

OpenStudy (anonymous):

sorry took me a bit to write it out, wanted to make sure I got it right :)

Mehek (mehek14):

put 2 for m \(\tt{y=2x+b}\) now choose one of the 2 ordered pairs

OpenStudy (anonymous):

6 @Mehek14